Can someone please help me with this? I have been trying to solve it and I am now on my last attempt and I don't know what IHome Read, Study & Practice Gradebook ORION Downloadable eTextbook Assignment> Open Assignment CALCULATOR FULL SCREEN PRINTER VERSION BACK NEXT ASSIGNMENT RESOURCES Ch. 6.Homework Your answer is incorrect. Try again Practice Exercise 6- What is the future value of $29,000 at the end of 5 periods at 11% compounded interest? (Round answer to 0 decimal places, eg. 458,581.) Future value 19,326 Review Score y Study Objective LINK TO TEXT LINK TO TEXT LINK TO TEXT Your answer is incorrect. Try again what is the present value of $29,000 to be received at the end of each of 20 periods, discounted at 8% compound interest? (Round answer to。 decimal places, e.g. 458,581.) Present value 127,128 Reflect in ePortfolioHome Read, Study & Practice Assignment Gradebook ORION Downloadable eTextbook Assignment>Open Assignment CALCULATOR FULL SCREEN PRINTER VERSION BACK NEXT ASSIGNMENT RESOURCES Ch.6.Homework ractice Exercise 6 x Your answer is incorrect. Try again. what is the present value of $29,000 due 8 penods hence, discounted at 14%? (Round answer to O decinnal places, eg. 458,581 Practice Exercise 6-4 Practice Exercise 6-5 Present value 4513 Review Score Review Results by Study Objective Your answer is incorrect. Try again what is the future value of 15 periodic payments of $29,000 each made at the end of each period and compounded at 13%2 (Round answer to 0 decimal places, e.g. 458,581.) Future value 447,3 Reflect in ePortfolio am doing wrong
Answer #1
1) Future Value $           48,867
Working:
Future Value = Investment *(1+i)^n Where,
= 29000*(1+0.11)^5 i 11%
= $           48,867 n 5
2) Present Value $       2,84,726
Working:
Cumulative discount factor = (1-(1+i)^-n)/i Where,
= (1-(1+0.08)^-20)/0.08 i 8%
= 9.81814741 n 20
Present Value = 29000 x 9.818147
= $       2,84,726
3) Present Value $           10,166
Working:
Present Value = Future Value *(1+i)^-n Where,
= 29000*(1+0.14)^-8 i 14%
= $           10,166 n 8
4) Future Value $     11,72,106
Working:
Future Value of annuity of 1 = (((1+i)^n)-1)/i Where,
= (((1+0.13)^15)-1)/0.13 i 13%
= 40.4174644 n 15
Future Value = $       29,000 x 40.41746
= $ 11,72,106